OpenStudy (anonymous):

Solve for x. |x – 2| + 10 = 12

OpenStudy (anonymous):

x=4

OpenStudy (lucaz):

subtract 10 from both sides, you get |x-2|=2 the absolute value of a positive or negative value is always a positive number then you can solve x-2=2 and x-2=-2

OpenStudy (anonymous):

to me this equation is mental math probaly because i am in a higher grade or sometHing but it Is easy as pie